500 ml of gaseous hydrocarbon when burnt in excess of of O2 gave 2.5 l of CO2 and 3.0 l of H20(vap) the hydrocarbon is

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Write the eqn. first and balance it.
   CxHy          +         (2x + y/2)O2            =======>                xCO2                   +               (y/2)H2O
 moles = 0.5/22.4                                                                       moles = 2.5/22.4                     moles = 3/22.4  Now,
from limiting reagent concept,
     0.0223/1 = 0.111/x
or
x = 4.9 approx. 5
thus, x = 5
and also
0.0233/1 = 0.1339*2/y
0.0233 = 0.2678/y
or
y =  11.5 that is approx. we woukd take 12.
therfore y =12
the hydrocarbon is:
C5H12.